You have $81,871 to invest in two stocks and the risk-free security. Stock A has an expected return of 13.17 percent and Stock B has an expected return of 10.03 percent. You want to own $29,241 of Stock B. The risk-free rate is 5.21 percent and the expected return on the market is 11.24 percent. If you want the portfolio to have an expected return equal to that of the market, how much should you invest (in $) in the risk-free security? Answer to two decimals. (Hint: A negative answer is OK - it means you borrowed (rather than lent or invested) at the risk free rate.) Thank you!!
Answer #1

Let x be invested in risk free security

29241*10.03% + x*5.21% + (81871-29241-x)*13.17% = 81871*11.24%

2932.8723 + 0.0521x + 6931.371 – 0.1317x = 9203.3004

X = $8,315.87

Hence, the answer is $8,315.87